# Narrowing the Focus: Learned Optimizers for Pretrained Models

### v2

@@ 6 Ablations and Variants
We first show the results of using either SGD or Adam alone as the wrapped optimizer rather than both together.
@@
Without providing layer embeddings to L3RS, the optimizer would have more difficulty identifying which layer it was operating on needing to rely heavily on the adaptive EMA features. An alternative to the layer embeddings is to meta-learn a separate MLP model for each layer of the target model. This strategy greatly increases the total parameter count of the optimizer. We investigate both the per-layer MLP style, as well as the single MLP without layer embeddings. See Table 2 for results.
@@
We additionally train models with different adaptive EMA feature smoothing factors. While L3RS uses $\gamma\in\{0.99,0.9,0.0\}$, we also test $\gamma\in\{0.9,0.0\}$, $\gamma\in\{0.0\}$ (which corresponds to the raw features without averaging), and $\gamma\in\varnothing$ (which corresponds to not receiving the adaptive features at all). See Table 3 for results.
@@ 7 Conclusions
Learned optimization is a powerful paradigm which has the potential to outperform existing methods. Our results suggest that narrowing the domain that learned optimizers are meta-trained on can provide a reduced meta-training cost, and allow learning domain specific exploitations that can increase their performance, especially on fine-tuning tasks. We provide results repurposing VeLO, an architecture designed for general learned optimization, as well as propose a new architecture L3RS, designed to take advantage of the narrow domain. We also demonstrate the robustness of this method on out of distribution tasks. As fine-tuning tasks continue to become more relevant, especially for LLMs, this paradigm and architecture provides a promising direction for future research. Improving this technique further can result in faster model training, better performance, and robust data generalization.
@@ Appendix B Meta-Training
Figures 7 and 7 show the example of the relative and absolute time features used by L3RS.
Appendix C Evaluation Tables
We provide the full evaluation tables from the main evaluation experiments Figure 5 A-E. Table 5 provides the results from the in-domain evaluation A. Table 6 provides the results for all head-only fine-tuning for evaluation A. Table 7 provides the results from the dataset-generalization evaluation B. Table 8 provides the results from the checkpoint-generalization evaluation C. Table 9 provides the results from the initialization and dataset-generalization evaluation D. Table 10 provides the results from the RandomInit evaluation E.

### v3

@@ 6 Ablations and Variants
We will show the results of using either SGD or Adam alone as the wrapped optimizer rather than both together.
@@
Without layer embeddings, L3RS struggles to distinguish between layers, relying heavily on adaptive EMA features. Alternatively, we can meta-learn a separate MLP for each layer of the target model. This significantly increases the optimizer’s parameter count. We compare this per-layer MLP approach with a single shared MLP (without layer embeddings). We additionally report results for a Global method, which uses a single learning rate ($\lambda$) and direction weights ($\mu_{p}$) for all layers. Results for these different optimizer configurations are presented in Table 2.
@@
We also investigated the impact of different smoothing factors ($\gamma$) for the adaptive exponential moving average (EMA) features. The standard L3RS uses $\gamma\in\{0.99,0.9,0.0\}$. For comparison, we trained models with $\gamma\in\{0.9,0.0\}$, $\gamma=0.0$ (representing raw features without averaging), and no adaptive EMA features ($\gamma=\emptyset$). The results are summarized in Table 3.
@@ 7 Conclusions
Learned optimization is a powerful paradigm which has the potential to outperform existing methods. Our results suggest that narrowing the domain that learned optimizers are meta-trained on can provide a reduced meta-training cost, and allow learning domain specific exploitations that can increase their performance, especially on fine-tuning tasks. We provide results repurposing VeLO, an architecture designed for general learned optimization, as well as propose a new architecture L3RS, designed to take advantage of the narrow domain. We demonstrate the robustness of this method on out of distribution tasks. As fine-tuning tasks continue to become more relevant, especially for LLMs, this paradigm and architecture provides a promising direction for future research. Improving this technique further can result in faster model training, better performance, and robust data generalization.
@@ Appendix B Meta-Training
Figures 8 and 8 show the example of the relative and absolute time features used by L3RS.
Appendix C Learned mixing coefficients dynamics
To better understand the optimization dynamics of L3RS during a 100-evaluation run, we visualize the learning rate and direction mix per layer. Figure 9 provides a more granular view, displaying the same relationship for each individual layer.
Appendix D Evaluation Tables
We provide the full evaluation tables of accuracy and loss from the main evaluation experiments Figure 5 A-E. We additionally reproduce Figure 5 using the evaluation losses in Figure 10. Table 5 and Table  6 provide the accuracy and loss results respectively from the in-domain evaluation A. Table 7 and Table 8 provide the accuracy and loss results respectively for all head-only fine-tuning for evaluation A. Table 9 and Table 10 provide the accuracy and loss results respectively from the dataset-generalization evaluation B. Table 11 and Table 12 provide the accuracy and loss results respectively from the checkpoint-generalization evaluation C. Table 13 and Table 14 provide the accuracy and loss results respectively from the initialization and dataset-generalization evaluation D. Table 15 and Table 16 provide the accuracy and loss results respectively from the RandomInit evaluation E.
@@ A $\widetilde{\theta}_{0}$: ImageNet,   ${\widetilde{D}_{K}^{t}},\widetilde{D}^{e}$: ImageNet25Eval
Refer to caption F
Refer to caption Refer to caption
Figure 10: Meta-evaluation of the L3RS, meta-trained on ImageNet25 for $10$ to $500$ steps along with various benchmarks. Performance is compared to VeLO (left), Adam with cosine learning rate (center), and Adam with a constant learning rate (right). Each marker represents model loss evaluation at that number of steps. Solid lines indicate the number of steps for in-distribution evaluation, while dashed lines indicate generalization to more steps than meta-training.
F. Speedup of L3RS. For in-domain generalization, this shows how much faster L3RS achieves a given loss compared to the baselines.
@@
L3RS $1.17\pm 0.13$ $0.82\pm 0.11$ $0.67\pm 0.10$ $0.56\pm 0.10$ $0.46\pm 0.09$ $0.39\pm 0.08$ $0.38\pm 0.09$ $0.43\pm 0.11$
VeLO (ft) $1.35\pm 0.17$ $0.90\pm 0.13$ $0.75\pm 0.11$ $0.62\pm 0.11$ $0.50\pm 0.09$ $0.45\pm 0.10$ $0.92\pm 0.96$ $3.23\pm 0.12$
VeLO (og) $1.75\pm 0.18$ $1.81\pm 0.16$ $1.57\pm 0.14$ $1.33\pm 0.14$ $0.98\pm 0.13$ $0.74\pm 0.12$ $0.61\pm 0.11$ $0.67\pm 0.14$
Adam (cosine), 1e-3 $1.67\pm 0.15$ $1.20\pm 0.14$ $1.01\pm 0.13$ $0.82\pm 0.13$ $0.61\pm 0.11$ $0.50\pm 0.10$ $0.45\pm 0.09$ $0.49\pm 0.13$
Adam (cosine), 1e-4 $3.18\pm 0.06$ $2.63\pm 0.09$ $1.94\pm 0.13$ $1.23\pm 0.14$ $0.69\pm 0.11$ $0.51\pm 0.10$ $0.42\pm 0.09$ $0.40\pm 0.10$
Adam (cosine), 1e-5 $3.56\pm 0.05$ $3.49\pm 0.05$ $3.38\pm 0.05$ $3.17\pm 0.06$ $2.60\pm 0.09$ $1.88\pm 0.13$ $1.17\pm 0.14$ $0.66\pm 0.11$
Adam (const), 1e-3 $1.44\pm 0.15$ $1.34\pm 0.14$ $1.20\pm 0.15$ $1.00\pm 0.13$ $0.83\pm 0.12$ $0.72\pm 0.12$ $0.66\pm 0.11$ $0.70\pm 0.15$
Adam (const), 1e-4 $2.80\pm 0.08$ $1.95\pm 0.13$ $1.24\pm 0.14$ $0.80\pm 0.12$ $0.56\pm 0.10$ $0.46\pm 0.09$ $0.43\pm 0.09$ $0.46\pm 0.12$
Adam (const), 1e-5 $3.51\pm 0.05$ $3.38\pm 0.05$ $3.17\pm 0.06$ $2.78\pm 0.08$ $1.89\pm 0.13$ $1.17\pm 0.14$ $0.74\pm 0.11$ $0.50\pm 0.10$
Table 6: Average and standard deviation evaluation loss for main experiment (A).
@@
VeLO (og) $2.21\pm 0.11$ $1.11\pm 0.14$ $0.92\pm 0.14$ $0.85\pm 0.13$ $0.76\pm 0.13$ $0.68\pm 0.12$ $0.65\pm 0.12$ $0.66\pm 0.12$
Adam (cosine), 1e-3 $3.20\pm 0.05$ $2.88\pm 0.07$ $2.45\pm 0.09$ $1.86\pm 0.11$ $1.21\pm 0.12$ $0.93\pm 0.12$ $0.77\pm 0.11$ $0.66\pm 0.11$
Adam (cosine), 1e-4 $3.55\pm 0.05$ $3.48\pm 0.05$ $3.38\pm 0.05$ $3.23\pm 0.05$ $2.89\pm 0.06$ $2.46\pm 0.08$ $1.88\pm 0.11$ $1.21\pm 0.12$
Adam (cosine), 1e-5 $3.60\pm 0.05$ $3.59\pm 0.05$ $3.58\pm 0.05$ $3.55\pm 0.05$ $3.48\pm 0.05$ $3.39\pm 0.05$ $3.23\pm 0.05$ $2.90\pm 0.06$
Adam (const), 1e-3 $3.00\pm 0.06$ $2.47\pm 0.09$ $1.87\pm 0.11$ $1.35\pm 0.12$ $0.95\pm 0.12$ $0.79\pm 0.11$ $0.70\pm 0.11$ $0.64\pm 0.11$
Adam (const), 1e-4 $3.50\pm 0.05$ $3.38\pm 0.05$ $3.23\pm 0.05$ $3.00\pm 0.06$ $2.47\pm 0.08$ $1.89\pm 0.11$ $1.35\pm 0.12$ $0.93\pm 0.12$
Adam (const), 1e-5 $3.59\pm 0.05$ $3.58\pm 0.05$ $3.55\pm 0.05$ $3.51\pm 0.05$ $3.39\pm 0.05$ $3.24\pm 0.05$ $3.00\pm 0.06$ $2.46\pm 0.08$
Table 8: Average and standard deviation evaluation loss for all Head-only fine-tuning from main experiment (A).
@@
L3RS $1.44\pm 0.10$ $1.02\pm 0.09$ $0.85\pm 0.09$ $0.73\pm 0.10$ $0.61\pm 0.09$ $0.55\pm 0.08$ $0.49\pm 0.08$ $0.45\pm 0.08$
VeLO (ft) $1.73\pm 0.25$ $1.13\pm 0.10$ $0.95\pm 0.09$ $0.82\pm 0.26$ $0.66\pm 0.09$ $0.78\pm 0.51$ $3.65\pm 15.11$ $1209.28\pm 11410.45$
VeLO (og) $1.88\pm 0.12$ $1.80\pm 0.12$ $1.57\pm 0.12$ $1.34\pm 0.10$ $1.01\pm 0.11$ $0.79\pm 0.10$ $0.64\pm 0.09$ $0.53\pm 0.08$
Adam (cosine), 1e-3 $1.82\pm 0.10$ $1.33\pm 0.09$ $1.11\pm 0.10$ $0.91\pm 0.10$ $0.71\pm 0.09$ $0.61\pm 0.08$ $0.54\pm 0.08$ $0.47\pm 0.08$
Adam (cosine), 1e-4 $3.19\pm 0.06$ $2.75\pm 0.06$ $2.16\pm 0.08$ $1.42\pm 0.09$ $0.85\pm 0.09$ $0.66\pm 0.09$ $0.55\pm 0.08$ $0.46\pm 0.08$
Adam (cosine), 1e-5 $3.52\pm 0.06$ $3.46\pm 0.06$ $3.37\pm 0.06$ $3.21\pm 0.06$ $2.75\pm 0.06$ $2.10\pm 0.09$ $1.35\pm 0.10$ $0.80\pm 0.09$
Adam (const), 1e-3 $1.61\pm 0.12$ $1.45\pm 0.11$ $1.26\pm 0.10$ $1.09\pm 0.11$ $0.90\pm 0.10$ $0.80\pm 0.10$ $0.71\pm 0.10$ $0.61\pm 0.09$
Adam (const), 1e-4 $2.91\pm 0.06$ $2.18\pm 0.08$ $1.46\pm 0.09$ $0.98\pm 0.10$ $0.72\pm 0.09$ $0.61\pm 0.08$ $0.55\pm 0.08$ $0.48\pm 0.08$
Adam (const), 1e-5 $3.48\pm 0.06$ $3.38\pm 0.06$ $3.21\pm 0.06$ $2.90\pm 0.06$ $2.10\pm 0.09$ $1.36\pm 0.10$ $0.90\pm 0.09$ $0.63\pm 0.09$
Table 10: Average and standard deviation evaluation loss for main experiment (B).
@@
L3RS $1.34\pm 0.13$ $0.95\pm 0.12$ $0.78\pm 0.11$ $0.66\pm 0.11$ $0.54\pm 0.10$ $0.47\pm 0.09$ $0.44\pm 0.09$ $0.50\pm 0.11$
VeLO (ft) $1.64\pm 0.27$ $1.02\pm 0.13$ $0.88\pm 0.13$ $0.73\pm 0.12$ $0.59\pm 0.10$ $0.52\pm 0.09$ $0.58\pm 0.40$ $15.24\pm 39.84$
VeLO (og) $1.66\pm 0.16$ $1.73\pm 0.15$ $1.55\pm 0.15$ $1.31\pm 0.14$ $0.97\pm 0.14$ $0.75\pm 0.12$ $0.62\pm 0.11$ $0.71\pm 0.15$
Adam (cosine), 1e-3 $1.75\pm 0.13$ $1.33\pm 0.15$ $1.14\pm 0.13$ $0.95\pm 0.13$ $0.72\pm 0.11$ $0.60\pm 0.10$ $0.52\pm 0.11$ $0.55\pm 0.13$
Adam (cosine), 1e-4 $3.05\pm 0.05$ $2.63\pm 0.07$ $2.05\pm 0.10$ $1.36\pm 0.13$ $0.79\pm 0.11$ $0.59\pm 0.10$ $0.48\pm 0.09$ $0.46\pm 0.10$
Adam (cosine), 1e-5 $3.35\pm 0.05$ $3.30\pm 0.05$ $3.22\pm 0.05$ $3.07\pm 0.05$ $2.63\pm 0.07$ $1.98\pm 0.11$ $1.28\pm 0.12$ $0.74\pm 0.11$
Adam (const), 1e-3 $1.60\pm 0.16$ $1.48\pm 0.14$ $1.33\pm 0.15$ $1.15\pm 0.13$ $0.94\pm 0.13$ $0.81\pm 0.12$ $0.74\pm 0.12$ $0.76\pm 0.16$
Adam (const), 1e-4 $2.79\pm 0.06$ $2.08\pm 0.10$ $1.39\pm 0.12$ $0.93\pm 0.12$ $0.65\pm 0.10$ $0.54\pm 0.09$ $0.49\pm 0.09$ $0.55\pm 0.12$
Adam (const), 1e-5 $3.32\pm 0.05$ $3.22\pm 0.05$ $3.07\pm 0.05$ $2.77\pm 0.06$ $1.99\pm 0.11$ $1.28\pm 0.12$ $0.83\pm 0.11$ $0.56\pm 0.10$
Table 12: Average and standard deviation evaluation loss for main experiment (C).
@@
L3RS $2.90\pm 0.08$ $2.72\pm 0.10$ $2.59\pm 0.10$ $2.41\pm 0.12$ $2.06\pm 0.14$ $1.79\pm 0.13$ $1.54\pm 0.14$ $1.35\pm 0.16$
VeLO (ft) $2.93\pm 0.13$ $2.82\pm 0.09$ $2.91\pm 0.48$ $14.01\pm 28.82$ $2552.78\pm 23833.64$ $14000.60\pm 138939.16$ $20165.58\pm 184680.17$ $88.27\pm 713.91$
VeLO (og) $2.85\pm 0.09$ $2.72\pm 0.10$ $2.60\pm 0.11$ $2.41\pm 0.13$ $1.95\pm 0.15$ $1.57\pm 0.14$ $1.22\pm 0.14$ $0.95\pm 0.15$
Adam (cosine), 1e-3 $2.98\pm 0.06$ $2.81\pm 0.08$ $2.67\pm 0.09$ $2.50\pm 0.11$ $2.13\pm 0.14$ $1.84\pm 0.14$ $1.52\pm 0.14$ $1.13\pm 0.13$
Adam (cosine), 1e-4 $3.26\pm 0.03$ $3.18\pm 0.03$ $3.08\pm 0.04$ $2.93\pm 0.06$ $2.68\pm 0.08$ $2.42\pm 0.11$ $2.00\pm 0.13$ $1.49\pm 0.13$
Adam (cosine), 1e-5 $3.31\pm 0.03$ $3.30\pm 0.03$ $3.29\pm 0.03$ $3.26\pm 0.03$ $3.18\pm 0.03$ $3.07\pm 0.04$ $2.91\pm 0.06$ $2.65\pm 0.09$
Adam (const), 1e-3 $2.88\pm 0.07$ $2.73\pm 0.09$ $2.61\pm 0.10$ $2.44\pm 0.12$ $2.14\pm 0.14$ $1.86\pm 0.15$ $1.56\pm 0.14$ $1.20\pm 0.15$
Adam (const), 1e-4 $3.21\pm 0.03$ $3.08\pm 0.04$ $2.93\pm 0.06$ $2.76\pm 0.08$ $2.46\pm 0.11$ $2.10\pm 0.12$ $1.76\pm 0.14$ $1.37\pm 0.14$
Adam (const), 1e-5 $3.31\pm 0.03$ $3.29\pm 0.03$ $3.26\pm 0.03$ $3.21\pm 0.03$ $3.07\pm 0.04$ $2.91\pm 0.06$ $2.72\pm 0.08$ $2.36\pm 0.11$
Table 14: Average and standard deviation evaluation loss for main experiment (D).
@@
L3RS $1.07\pm 0.09$ $0.76\pm 0.08$ $0.65\pm 0.09$ $0.58\pm 0.09$ $0.51\pm 0.08$ $0.47\pm 0.08$ $0.43\pm 0.07$ $0.40\pm 0.08$
VeLO (ft) $1.22\pm 0.25$ $0.77\pm 0.09$ $0.69\pm 0.09$ $0.60\pm 0.09$ $0.52\pm 0.08$ $1.29\pm 3.69$ $3.53\pm 12.36$ $24563.19\pm 163714.77$
VeLO (og) $1.38\pm 0.12$ $1.49\pm 0.12$ $1.33\pm 0.11$ $1.14\pm 0.11$ $0.85\pm 0.10$ $0.69\pm 0.09$ $0.57\pm 0.09$ $0.49\pm 0.08$
Adam (cosine), 1e-3 $1.52\pm 0.10$ $1.11\pm 0.08$ $0.97\pm 0.10$ $0.82\pm 0.10$ $0.67\pm 0.09$ $0.58\pm 0.09$ $0.52\pm 0.08$ $0.46\pm 0.08$
Adam (cosine), 1e-4 $2.98\pm 0.06$ $2.50\pm 0.07$ $1.85\pm 0.08$ $1.12\pm 0.09$ $0.67\pm 0.08$ $0.54\pm 0.08$ $0.47\pm 0.08$ $0.41\pm 0.07$
Adam (cosine), 1e-5 $3.34\pm 0.06$ $3.28\pm 0.06$ $3.19\pm 0.06$ $3.01\pm 0.06$ $2.50\pm 0.07$ $1.78\pm 0.08$ $1.05\pm 0.08$ $0.63\pm 0.08$
Adam (const), 1e-3 $1.38\pm 0.12$ $1.30\pm 0.11$ $1.17\pm 0.10$ $1.03\pm 0.11$ $0.88\pm 0.11$ $0.79\pm 0.09$ $0.70\pm 0.09$ $0.61\pm 0.10$
Adam (const), 1e-4 $2.69\pm 0.07$ $1.89\pm 0.08$ $1.15\pm 0.09$ $0.77\pm 0.08$ $0.60\pm 0.08$ $0.53\pm 0.08$ $0.49\pm 0.08$ $0.45\pm 0.08$
Adam (const), 1e-5 $3.30\pm 0.06$ $3.19\pm 0.06$ $3.01\pm 0.06$ $2.67\pm 0.07$ $1.78\pm 0.08$ $1.05\pm 0.08$ $0.69\pm 0.08$ $0.51\pm 0.08$
Table 16: Average and standard deviation evaluation loss for main experiment (E).